# Unleashing Reasoning Capability of LLMs via Scalable Question Synthesis from Scratch

### v1

@@ Abstract
The availability of high-quality data is one of the most important factors in improving the reasoning capability of LLMs. Existing works have demonstrated the effectiveness of creating more instruction data from seed questions or knowledge bases. Recent research indicates that continually scaling up data synthesis from strong models (e.g., GPT-4) can further elicit reasoning performance. Though promising, the open-sourced community still lacks high-quality data at scale and scalable data synthesis methods with affordable costs. To address this, we introduce ScaleQuest, a scalable and novel data synthesis method that utilizes “small-size” (e.g., 7B) open-source models to generate questions from scratch without the need for seed data with complex augmentation constraints. With the efficient ScaleQuest, we automatically constructed a mathematical reasoning dataset consisting of 1 million problem-solution pairs, which are more effective than existing open-sourced datasets. It can universally increase the performance of mainstream open-source models (i.e., Mistral, Llama3, DeepSeekMath, and Qwen2-Math) by achieving 29.2% to 46.4% gains on MATH. Notably, simply fine-tuning the Qwen2-Math-7B-Base model with our dataset can even surpass Qwen2-Math-7B-Instruct, a strong and well-aligned model on closed-source data, and proprietary models such as GPT-4-Turbo and Claude-3.5 Sonnet. [^1] Code, data, and models are publicly available: https://github.com/yyDing1/ScaleQuest.
Figure 1: Left: Results of different models on MATH, where -ScaleQuest denotes ours. Right: Results of Llama3-8B fine-tuned on publicly available datasets constructed by different methods.
@@ 1 Introduction
How to improve the reasoning capabilities of Large Language Models (LLMs) has attracted significant attention. The success of recent advanced models, such as OpenAI o1 and Claude-3.5, heavily depends on access to extensive, diverse, and high-quality reasoning datasets. However, the proprietary nature of the data presents a significant barrier to the open-source community. Recent works have highlighted data synthesis as a promising approach (Ntoutsi et al., 2020) to address data scarcity for instruction tuning (Inan et al., 2023). As recent works have disclosed that crafting the right questions is crucial for eliciting the reasoning capabilities of LLMs (Yu et al., 2023a; Shah et al., 2024), the core of reasoning data synthesis lies in creating large-scale and novel questions.
Previous efforts in reasoning data synthesis have demonstrated the effectiveness of leveraging powerful language models to generate instructions. We categorize these approaches into two types: question-driven approaches and knowledge-driven approaches. Question-driven methods include question rephrasing (Yu et al., 2023a), evol-instruct (Xu et al., 2023; Luo et al., 2023; Zeng et al., 2024), question back-translation (Lu et al., 2024), or providing few-shot examples (Mitra et al., 2024). These methods are limited in data diversity, as the generated problems closely resemble the seed questions, with only minor modifications such as added conditions or numerical changes. This lack of diversity hampers their scalability potential. To improve question diversity, recent knowledge-driven works (Huang et al., 2024b) scale question synthesis by constructing knowledge bases (Li et al., 2024b) or concept graphs (Tang et al., 2024) and sampling key points (Huang et al., 2024a) from them to generate new questions. Nevertheless, the above two types of approaches commonly rely on strong models, like GPT-4, to synthesize new questions, but the high API costs make it impractical to generate large-scale data. As a result, despite these advancements, the open-source community still faces a shortage of high-quality data at scale and cost-effective synthesis methods.
@@
We generated data based on two lightweight, open-source models: DeepSeekMath-7B-RL (Shao et al., 2024) and Qwen2-Math-7B-Instruct (Yang et al., 2024a), producing a final dataset of 1 million question-answer pairs. As shown in Figure 1, our synthetic dataset boosts performance by 29.2% to 46.4% across four major open-source models: Mistral-7B (Jiang et al., 2023), Llama3-8B (Dubey et al., 2024), DeepSeekMath-7B (Shao et al., 2024), and Qwen2-Math-7B (Yang et al., 2024a). Compared with other publicly available datasets such as MetaMath (Yu et al., 2023a), DART-Math (Tong et al., 2024), and NuminaMath (Li et al., 2024c), our approach demonstrates great scalability in both in-domain and out-of-domain evaluation. In terms of in-domain evaluation, our method outperforms existing high-quality open-source datasets, achieving better results with the same amount of data. For out-of-domain evaluation, compared with other datasets, the performance of our synthetic dataset continues to show promising trends as the volume of training data increases, indicating significant potential for further improvements through ongoing data scaling.
@@ 2 ScaleQuest: Scaling Question Synthesis from Scratch
In this section, we first explain the motivation and process of our question generation method (section 2.1). Then, we introduce how to train a question generator via Question Fine-Tuning (section 2.2) and Question Preference Optimization (section 2.3). Next, we use the question generator to generate math questions, followed by a filtering process (section 2.4). Finally, we describe the response generation process (section 2.5). The overview of our method is illustrated in Figure 2.
@@ 2.2 Question Fine-Tuning (QFT)
To activate the model’s question generation capability, we first perform Question Fine-Tuning (QFT), where we train the problem-solving model using a small set of problems. To ensure that the generator stops after producing the questions and does not continue generating a response, we added an end-of-sentence token at the end of each question. We used approximately 15K problems (without solutions) by mixing the training set of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) datasets as training samples. We train DeepSeekMath-7B-RL Shao et al. (2024) and Qwen2-Math-7B-Instruct Yang et al. (2024a) with these samples.
The purpose of utilizing these problems is to activate the model’s question-generation capability rather than to make the model memorize them. To validate this hypothesis, we trained the model separately using the GSM8K and MATH datasets and compared whether the distribution of the generated questions matched that of the training data. To evaluate the question distribution, we used a difficulty classifier, which maps a question into a difficulty score (details in Section 2.4). We performed QFT based on Qwen2-Math-7B (Yang et al., 2024a), then used the two QFT models, Qwen2-QFT-GSM8K and Qwen2-QFT-MATH, to synthesize 10K questions. The difficulty distribution of these four datasets is shown in Figure 3. We found that the generated questions separately differed from both GSM8K and MATH, yet they both converged toward the same distribution. Additionally, the QFT model, trained on English questions, demonstrated the ability to generate a substantial number of questions in other languages. Both phenomena suggest that the QFT process enhances the model’s question-generation capabilities without leading to overfitting the training data.
@@ 2.3 Question Preference Optimization (QPO)
Figure 4: The solvability and difficulty of the raw questions generated by the QFT model and the optimized ones.
The model is able to generate meaningful and diverse questions after QFT, but the quality is still not high enough, as shown in Figure 2. This is reflected in two aspects: (1) solvability: the math problem should have appropriate constraints and correct answers, and (2) difficulty: the model needs to learn from more challenging problems, yet some of the generated questions are still too simple. To address these two aspects, we applied Question Preference Optimization (QPO).
We first used the model after QFT to generate 10K questions. Then, we optimized these samples using an external LLM, focusing primarily on solvability and difficulty. We found that simultaneously optimizing both posed a challenge for the LLMs. Therefore, for each sample, we randomly selected one of the two optimization directions, prioritizing either solvability or difficulty. The optimization prompts can be found in Figure 9 and 10. The optimized questions, denoted as $y_{w}$, are treated as preferred data, while the original questions before optimization, denoted as $y_{l}$, are considered dispreferred data. We modified the loss for Direct Preference Optimization (DPO) (Rafailov et al., 2024) formulation to fit our approach:
$\mathcal{L}_{\text{QPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w})}{\pi_{\text{ref}}(y_{w})}-\beta\log\frac{\pi_{\theta}(y_{l})}{\pi_{\text{ref}}(y_{l})}\right)\right].$ (2)
The question optimization process placed significant demands on the model’s ability to follow complex instructions. We experimented with two question optimization models: Qwen2-Math-7B-Instruct and GPT-4o-mini. To evaluate improvements in solvability and difficulty, we used GPT-4o, with the prompts for this evaluation provided in Figure 11 and 12. The results are shown in Figure 4. In terms of solvability, Qwen2-Math-7B-Instruct proved inadequate for this task, as the optimized questions resulted in decreased solvability. A possible reason for this is the model’s insufficient ability to follow instructions accurately, resulting in many answers that fail to meet the specified optimization constraints. Consequently, we selected GPT-4o-mini as the question optimization model.
@@ 2.4 Question Filtering
After the QFT and QPO phases, we obtained two question generators: DeepSeekMath-QGen and Qwen2-Math-QGen. There are still some minor issues in the generated questions, primarily related to language, solvability, and difficulty. To address these challenges, we applied the following filtering steps:
@@
The question generator models still produce a substantial number of math questions in other languages, accounting for approximately 20%. Since our focus is on English math questions, we removed non-English questions by identifying questions containing non-English characters and filtering out those samples.
Solvability Filering
Although QPO effectively enhances the solvability of generated questions, some questions remain nonsensical. This is primarily due to (1) poorly constrained questions, where missing conditions, redundant conditions, or logical inconsistencies occur, and (2) questions that do not yield meaningful outcomes (e.g., answers involving the number of people should result in a non-negative integer). To filter out such samples, we used Qwen2-Math-7B-Instruct to evaluate whether the question is meaningful and whether the conditions are sufficient. The prompts used for the solvability check are provided in Figure 11.
@@
We measure the difficulty of a question using the fail rate (Tong et al., 2024) — the proportion of incorrect responses when sampling $n$ responses for a given question. This metric aligns with the intuition that harder questions tend to result in fewer correct responses. Following Tong et al. (2024), we used DeepseekMath-7B-RL as the sampling model to evaluate the difficulty of each question in the training sets of GSM8K and MATH, obtaining the fail rate for each question as its difficulty score. We then used this data to train a difficulty scorer. Specifically, we built upon DeepseekMath-7B-Base and added a classification head on top of the model’s hidden state. The difficulty score $d$ is computed and optimized as:
$d=Wh_{l}+b,\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-d_{i})^{2},$ (3)
where $W$ and $b$ are the weights and biases of the classification head, $h_{l}$ represents the last hidden state of the sequence, and $d_{i}$ is the predicted difficulty score for the $i$-th question. The loss function $\mathcal{L}$ is the mean squared error (MSE), where $y_{i}$ represents the true difficulty score for the $i$-th question. We then used the scorer to predict the difficulty of each synthetic question and sample based on the question’s difficulty. Specifically, we filtered out a portion of the questions generated by DeepSeekMath-QGen that were overly simple. In contrast, the difficulty distribution of Qwen2-Math-QGen was more balanced, so no sampling was necessary.
2.5 Response Generation with Reward Filtering
Prior efforts to guarantee the quality of solutions include two aspects: (1) rejection sampling (Yuan et al., 2023): Large language models (LLMs) are tasked with generating multiple responses, specifically reasoning paths, for each instruction. Only reasoning paths that lead to the correct answer are preserved as solutions (Tong et al., 2024). (2) If the correct answer is unavailable, a majority voting method is used (Huang et al., 2024a), selecting the answer that appears most frequently across multiple reasoning paths and retaining these as the solutions. We use the reward model score as a metric for evaluating the quality of responses, considering its broader applicability, as there is often no single correct answer in other reasoning tasks like code generation and tool planning. Specifically, for each question, we generate 5 solutions and select the solution with the highest reward model scores as the preferred solution. In our experiments, we use InternLM2-7B-Reward (Cai et al., 2024) as our reward model.
@@ 3 Experiment
Table 1: Main results on four mathematical reasoning benchmarks. Bold means the best score within the respective base model. The baselines use different synthesis models, such as GPT-4, GPT-4-Turbo, GPT-4o, DeepSeekMath, and Qwen2-Math. If multiple models are used, only the latest released one is marked. More details concerning these datasets are shown in Figure 5.
@@ 3.1 Experimental Setup
Our question synthesis process relies on two problem designer models: Deepseek-QGen and Qwen2-Math-QGen, which were trained using QFT (section 2.2) and QPO (section 2.3), based on DeepSeekMath-7B-RL (Shao et al., 2024) and Qwen2-Math-7B-Instruct (Yang et al., 2024a), respectively. During the QFT stage, both models are trained on a mixed training subset of GSM8K and MATH problems, containing a total of 15K problems. We trained for only 1 epoch, considering that training for more epochs might cause the models to overfit the training problems and negatively impact the diversity of generated questions. We also used sequence packing (Krell et al., 2021) to accelerate training. In the QPO stage, we use 10K preference data for training, with a learning rate of 5e-7 and a batch size of 128.
Question Generation
The two question generation models were then utilized to generate a total of 2 million questions, with 1 million from each model. During this process, we set the maximum generation length to 512, a temperature of 1.0, and a top-p value of 0.99. To ensure quality, we applied a question filtering pipeline (section 2.4) that involved language filtering, solvability filtering, and difficulty sampling. This process refined the dataset, leaving approximately 1M questions to form the final question pool, 400K from Deepseek-QGen and 600K from Qwen2-Math-QGen.
Response Generation
Based on the problems, we synthesized responses (section 2.5) using Qwen2-Math-7B-Instruct (Yang et al., 2024a). In the process, we set the maximum generation length to 2048, with a temperature of 0.7 and top-p of 0.95. We use chain-of-thought prompt (Wei et al., 2022) to synthesize solutions. We use vLLM (Kwon et al., 2023) to accelerate the generation and Ray (Moritz et al., 2018) to deploy distributed inference. For each problem, we sampled 5 solutions and selected the one with the highest reward score as the final response. The final dataset consists of 1 million problem-solution pairs.
@@
We conducted instruction tuning on the synthetic problems and solutions using two general base models, Mistral-7B (Jiang et al., 2023) and Llama3-8B (Dubey et al., 2024), as well as two math-specialized base models, DeepSeekMath-7B (Shao et al., 2024) and Qwen2-Math-7B (Yang et al., 2024a). All models were fine-tuned for 3 epochs in our experiments unless specified otherwise. We used a linear learning rate schedule with a 3% warm-up ratio, reaching a peak of 5e-5 for Llama3 and DeepSeekMath and 1e-5 for the other models, followed by cosine decay to zero.
@@
We assessed the fine-tuned models’ performance across four datasets of increasing difficulty. Along with the widely used GSM8K (elementary level) and MATH (competition level), we included two more challenging benchmarks: College Math (Yuan et al., 2023) (college level) and Olympiad Bench (He et al., 2024) (Olympiad level). For evaluation, we employed the script from Tong et al. (2024) to extract final answers and determine correctness by comparing answer equivalency. The generated outputs were all in the form of natural language Chain-of-Thought (CoT) reasoning (Wei et al., 2022) through greedy decoding, with no tool integration, and we report zero-shot pass@1 accuracy.
@@
The main point of comparison is data synthesis methods, including: (1) WizardMath (Luo et al., 2023) proposes a reinforced Evol Instruct method; (2) MetaMath (Yu et al., 2023a) introduces three types of question bootstrapping; (3) MMIQC (Liu & Yao, 2024) proposes an iterative question composing method; (4) Orca-Math (Mitra et al., 2024) augments existing datasets using an Agent-Instruct method; (5) KPMath (Huang et al., 2024a) utilizes inherent topics and key points to synthesize problems; and (6) MathScale (Tang et al., 2024) builds a concept graph to generate new questions. In addition to this, we also involved other large math corpus like (7) DART-Math (Tong et al., 2024) enhances the response generation process through difficulty-guided rejection sampling; (8) Numina-Math (Li et al., 2024c) collects a large corpus by combining existing synthetic data with real-world datasets. More details of these datasets are shown in Table 5. We found that different scripts yielded varying evaluation results. To ensure consistency, we evaluated all released models using the same evaluation scripts. For methods without available results or released models, we retrained the models using their publicly available data.

### v2

# Unleashing LLM Reasoning Capability via Scalable Question Synthesis from Scratch

@@ Abstract
Improving the mathematical reasoning capabilities of Large Language Models (LLMs) is critical for advancing artificial intelligence. However, access to extensive, diverse, and high-quality reasoning datasets remains a significant challenge, particularly for the open-source community. In this paper, we propose ScaleQuest, a novel, scalable, and cost-effective data synthesis method that enables the generation of large-scale mathematical reasoning datasets using lightweight 7B-scale models. ScaleQuest introduces a two-stage question-tuning process comprising Question Fine-Tuning (QFT) and Question Preference Optimization (QPO) to unlock the question generation capabilities of problem-solving models. By generating diverse questions from scratch – without relying on powerful proprietary models or seed data – we produce a dataset of 1 million problem-solution pairs. Our experiments demonstrate that models trained on our data outperform existing open-source datasets in both in-domain and out-of-domain evaluations. Furthermore, our approach shows continued performance improvement as the volume of training data increases, highlighting its potential for ongoing data scaling. The extensive improvements observed in code reasoning tasks demonstrate the generalization capabilities of our proposed method. Our work provides the open-source community with a practical solution to enhance the mathematical reasoning abilities of LLMs. [^1] Project Page: https://scalequest.github.io.
Unleashing LLM Reasoning Capability via Scalable Question Synthesis from Scratch
Yuyang Ding1, Xinyu Shi1, Xiaobo Liang1, Juntao Li1thanks:  Corresponding author Zhaopeng Tu2, Qiaoming Zhu1, Min Zhang1 1Soochow University   2Tencent AI Lab {yyding23,xyshi02}@stu.suda.edu.cn {xbliang,ljt,qmzhu,minzhang}@suda.edu.cnzptu@tencent.com
@@ 1 Introduction
How to improve the mathematical reasoning capabilities of Large Language Models (LLMs) has attracted significant attention. The success of recent advanced models, such as OpenAI o1 and Claude-3.5, heavily depends on access to extensive, diverse, and high-quality reasoning datasets. However, the proprietary nature of the data presents a significant barrier to the open-source community. Recent works have highlighted data synthesis as a promising approach (Ntoutsi et al., 2020) to address data scarcity for instruction tuning (Inan et al., 2023). As recent works have disclosed that crafting the right questions is crucial for eliciting the reasoning capabilities of LLMs (Yu et al., 2023a; Shah et al., 2024), the core of reasoning data synthesis lies in creating large-scale and novel questions.
Figure 1: Results of Llama3-8B-Base fine-tuned on publicly available datasets in MATH and OlympiadBench. Our approach demonstrates strong scalability and significant potential for further improvement.
Previous efforts in reasoning data synthesis have demonstrated the effectiveness of leveraging powerful language models to generate instructions. We categorize these approaches into two types: question-driven approaches and knowledge-driven approaches. Question-driven methods include question rephrasing (Yu et al., 2023a), evol-instruct (Xu et al., 2023; Luo et al., 2023; Zeng et al., 2024), question back-translation (Lu et al., 2024), or providing few-shot examples (Mitra et al., 2024). These methods are limited in data diversity, as the generated problems closely resemble the seed questions, with only minor modifications, such as added conditions or numerical changes. This lack of diversity hampers their potential for scalability. To improve question diversity, recent knowledge-driven works (Huang et al., 2024b) scale question synthesis by constructing knowledge bases (Li et al., 2024b) or concept graphs (Tang et al., 2024) and sampling key points (Huang et al., 2024a) from them to generate new questions. Nevertheless, the above two types of approaches rely on strong models, like GPT-4, to synthesize new questions, but the high API costs make it impractical to generate large-scale data. Despite these advancements, the open-source community still lacks high-quality data at scale and cost-effective synthesis methods.
@@
We generate data using lightweight 7B-scale models, producing a final dataset of 1 million question-answer pairs. As shown in Figure 1, compared with other publicly available datasets such as MetaMath (Yu et al., 2023a), DART-Math (Tong et al., 2024), and NuminaMath (Li et al., 2024c), our approach demonstrates great scalability in both in-domain and out-of-domain evaluation. In terms of in-domain evaluation, our method outperforms existing high-quality open-source datasets, achieving better results with the same amount of data. For out-of-domain evaluation, the performance of our approach continues to show promising trends as the volume of training data increases, indicating significant potential for further improvements through ongoing data scaling. Additionally, we also extend our approach to long chain-of-thought and code reasoning tasks, demonstrating its effectiveness, with details in Appendix C.
@@ 2.2 Question Fine-Tuning (QFT)
To activate the model’s question generation capability, we first perform Question Fine-Tuning (QFT), where we train the problem-solving model using a small set of problems. To ensure that the generator stops after producing the questions and does not continue generating a response, we add an end-of-sentence token at the end of each question. We use approximately 15K problems (without solutions) by mixing the training set of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) datasets as training samples.
The purpose of utilizing these problems is to activate the model’s question-generation capability rather than to make the model memorize them. To validate this hypothesis, we train the model separately using the GSM8K and MATH datasets and compare whether the distribution of the generated questions matched that of the training data. To evaluate the question distribution, we use a difficulty classifier, which maps a question into a difficulty score (details in Section 2.4). We perform QFT based on Qwen2-Math-7B-Instruct (Yang et al., 2024a), then use the two QFT models, Qwen2-QFT-GSM8K and Qwen2-QFT-MATH, to synthesize 10K questions. The difficulty distribution of these four datasets is shown in Figure 3. We find that the generated questions separately differed from both GSM8K and MATH, yet they both converged toward the same distribution. This suggests that the QFT process enhances the model’s question-generation capabilities without leading to overfitting the training data.
@@ 2.3 Question Preference Optimization (QPO)
The model is able to generate meaningful and diverse questions after QFT, but the quality is still not high enough, as shown in Figure 2. This is reflected in two aspects: (1) solvability: the math problem should have appropriate constraints and correct answers, and (2) difficulty: the model needs to learn from more challenging problems, yet some of the generated questions are still too simple. To address these two aspects, we apply Question Preference Optimization (QPO).
We first use the model after QFT to generate 10K questions. Subsequently, we aim to refine these samples with a primary focus on solvability and difficulty. We leverage an external LLM for optimization as an alternative to manual annotation. However, we find that simultaneously optimizing both poses a challenge for the LLMs. Therefore, for each sample, we randomly select one of the two optimization directions, prioritizing either solvability or difficulty (with optimization prompts in Figure 10 and 11). The optimized questions, denoted as $y_{w}$, are treated as preferred data, while the original questions, denoted as $y_{l}$, are considered dispreferred data. Inspired by Direct Preference Optimization (DPO) (Rafailov et al., 2024), we propose QPO for question optimization:
Figure 4: The solvability and difficulty of the raw questions generated by the QFT model and the optimized ones. We use GPT-4o as a manual substitute to evaluate the effectiveness of solvability and difficulty optimization, with evaluation prompt in Figure 12 and 13.
$\mathcal{L}_{\text{QPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(y_{w},y_{l})\sim\mathcal{D}}\bigg{[}$ (2)
$\hskip 8.5359pt\log\sigma\bigg{(}\beta\log\frac{\pi_{\theta}(y_{w})}{\pi_{\text{ref}}(y_{w})}-\beta\log\frac{\pi_{\theta}(y_{l})}{\pi_{\text{ref}}(y_{l})}\bigg{)}\bigg{]}.$
In terms of model selection, we experiment with three models: Qwen2-Math-7B-Instruct, GPT-4o-mini, and Llama3.1-70B-Instruct. The results are shown in Figure 4. In terms of solvability, Qwen2-Math-7B-Instruct proves inadequate for this task, as the optimized questions result in decreased solvability. A possible reason for this is the model’s insufficient ability to follow instructions accurately, resulting in many answers that fail to meet the specified optimization constraints.
@@ 2.4 Question Filtering
After the QFT and QPO phases, we obtain the question generators Qwen2-Math-QGen. There are still some minor issues in the generated questions, primarily related to language, solvability, and difficulty. To address these challenges, we apply the filtering steps:
@@
The question generator models still produce a substantial number of math questions in other languages, accounting for approximately 20%. Since our focus is on English math questions, we remove non-English questions by identifying questions containing non-English characters and filtering out those samples.
Solvability Filtering
Although QPO effectively enhances the solvability of generated questions, some questions remain nonsensical. To filter out such samples, we use Qwen2-Math-7B-Instruct to evaluate whether the question is meaningful and whether the conditions are sufficient, with filtering prompts provided in Figure 12.
@@
We measure the difficulty of a question using the fail rate (Tong et al., 2024) — the proportion of incorrect responses when sampling $n$ responses for a given question. This metric aligns with the intuition that harder questions tend to result in fewer correct responses. Following Tong et al. (2024), we use DeepSeekMath-7B-RL as the sampling model to evaluate the difficulty of each question in the training sets of GSM8K and MATH, obtaining the fail rate for each question as its difficulty score. We then use this data to train a difficulty scorer based on DeepSeekMath-7B-Base with mean squared error (MSE) loss. We then use the scorer to predict the difficulty of each synthetic question. A portion of overly simple questions is then filtered out, which is empirically determined based on the difficulty distribution of the questions.
2.5 Response Generation
We use the reward model score as a metric for evaluating the quality of responses. For each question, we generate 5 solutions and select the solution with the highest reward model scores as the preferred solution. In our experiments, we use InternLM2-7B-Reward (Cai et al., 2024) as our reward model. This choice is primarily guided by the model’s performance on the reasoning subset of the Reward Bench (Lambert et al., 2024).
@@ 3 Experiment
Table 1: Main results on four mathematical reasoning benchmarks. Bold means the best score with the same base model. The baselines use different synthesis models for both question synthesis and response generation, such as GPT-3.5, GPT-4, and GPT-4o. For our approach, DSMath-7B-QGen and Qwen2-Math-7B-QGen are utilized for question synthesis, with Qwen2-Math-7B-Instruct used for response generation. If multiple models are used, only the latest released one is marked. More details about these datasets are shown in Table 7.
@@ 3.1 Experimental Setup
In addition to Qwen2-Math-QGen, we also train another question generator based on DeepSeekMath-7B-RL. We find that combining questions synthesized by multiple generators enhances data diversity, leading to improved final performance, as further discussed in section 3.3. In this process of the overall data synthesis, many models are mentioned in the context of QFT and QPO. We explain the inherent insights behind the model selection in Appendix D.
Question & Response Generation
The two question generation models are then utilized to generate a total of 2 million questions, with 1 million from each model. For difficulty filtering, we filter out a portion of overly simple questions generated by DeepSeekMath-QGen and keep the questions generated by Qwen2-Math-QGen as the difficulty distribution is more balanced. Based on the problems, we synthesize responses (section 2.5) using Qwen2-Math-7B-Instruct (Yang et al., 2024a). For each problem, we sample 5 solutions and select the one with the highest reward score as the final response. The final dataset consists of 1 million problem-solution pairs. A detailed analysis of our constructed dataset is provided in Appendix B.
@@
We conduct instruction tuning on the synthetic problems and solutions using two general base models, Mistral-7B (Jiang et al., 2023) and Llama3-8B (Dubey et al., 2024), as well as two math-specialized base models, DeepSeekMath-7B (Shao et al., 2024) and Qwen2-Math-7B (Yang et al., 2024a). More hyperparameters in the process are provided in Appendix A.
@@
We assess the fine-tuned models’ performance across four datasets of increasing difficulty. Along with the widely used GSM8K (elementary level) and MATH (competition level), we include two more challenging benchmarks: College Math (Tang et al., 2024) (college level) and Olympiad Bench (He et al., 2024) (Olympiad level). For evaluation, we employ the script from Tong et al. (2024) to extract final answers and determine correctness by comparing answer equivalency. The generated outputs are all in the form of natural language Chain-of-Thought (CoT) reasoning (Wei et al., 2022) through greedy decoding, with no tool integration, and we report zero-shot pass@1 accuracy. We also conduct more comprehensive evaluation, with details shown in Appendix E. We also extend our approach to the long cot and code reasoning task, with details and experiments provided in Appendix C.
@@
We mainly compare with data synthesis methods, including question-driven approach such as WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2023a), MMIQC (Liu and Yao, 2024), Orca-Math (Mitra et al., 2024) and knowledge-driven approach such as KPMath (Huang et al., 2024a), MathScale (Tang et al., 2024). In addition to this, we also involve other large math corpus like DART-Math (Tong et al., 2024) and Numina-Math (Li et al., 2024c). More details of these datasets are shown in Table 7.